OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. B(6, 2), m = - 1/2

OpenStudy (anonymous):

x+2y=10

OpenStudy (anonymous):

Y-y=m(X-x) put the point and slope into this equation Y-2=-1/2(X-6) Y-2=-1/2X+3 Y=-1/2X+5 can be written as previously multiply by 2 x+2y=10